# Enhancing Repository-Level Code Generation with Integrated Contextual Information

### v1

@@ 8. Conclusion and Future Work
In this paper, we aim to address the lack of type dependency information for statically typed programming languages in repository-level code generation. We propose CatCoder, a code generation framework for statically typed programming languages, which improves the performance of LLMs by retrieving relevant code and extracting type context using static analyzers. CatCoder is systematically evaluated on Java and Rust benchmarks that are built from real-world open-source repositories. The results show that CatCoder outperforms the RepoCoder baseline by up to 17.35%. In addition, we evaluate CatCoder with various LLMs. CatCoder improves the performance of all selected LLMs and therefore is generalizable. In future work, possible research attempts include investigating the data leakage problem of the benchmarks and improving the efficiency of CatCoder. These investigations may help further improve the applicability of our proposed approach in practical scenarios.

### v2

@@ 8. Conclusion and Future Work
In this paper, we aim to address the lack of type dependency information in repository-level code generation for statically typed programming languages. We propose CatCoder, a code generation framework for statically typed programming languages, which improves the performance of LLMs by retrieving relevant code and extracting type context using static analyzers. CatCoder is systematically evaluated on Java and Rust benchmarks that are built from real-world open-source repositories. The results show that CatCoder outperforms the RepoCoder baseline by up to 17.35% in terms of pass@k scores. Further studies demonstrate the generalizability and scalability of CatCoder. We release the replication package of CatCoder, including the source code of CatCoder, the benchmark datasets, and the evaluation scripts. In future work, possible research attempts include investigating the data leakage problem of the benchmarks. These investigations may help further improve the applicability of our proposed approach in practical scenarios.
Acknowledgements.
This research is supported by the Fundamental Research Funds for the Central Universities (No. 226-2025-00171).